OpenStudy (anonymous):

Find the y-intercept for this equation: 2x - 2y =8 -8 8 4 -4 -1 1

OpenStudy (e.mccormick):

Get it into y=mx+b form and b will be the y intercept.

OpenStudy (gorv):

add 2y on both side

OpenStudy (anonymous):

set x = 0 and solve for y. in std form it's very simple... eliminate the x term and solve what remains for y. that will be the y intercept.

OpenStudy (anonymous):

-2y = 8 what is y?

OpenStudy (gorv):

@pgpilot326 approch is best

OpenStudy (studygurl14):

2x - 2y = 8 Convert to this form: y = mx + b, where b = y-intercept -2y = -2x + 8 y = x - 4 y-intercept is -4

OpenStudy (gorv):

thn u dont need to convert anything

OpenStudy (gorv):

welcm to os :)

OpenStudy (gorv):

at y intercept x=0

OpenStudy (anonymous):

Find the y-intercept for this equation: \(2x - 2y =8 \) solve \[-2y=8\]in one step